OpenStudy (loser66):

Show that the ring M of n x n matrices over \(\mathbb R\) has no ideals other than 0 an M. Conclude that any ring homomorphism \(\phi:M\rightarrow S\) is either identically zero or is injective. Please, help

OpenStudy (loser66):

@misty1212

OpenStudy (anonymous):

the second statement is a direct consequence of the first i am not sure what you get to use the prove the first statement, but the following is true: if \(R\) is a ring and \(I\) an ideal in \(R\) then \(M_n(I)\) is an ideal in \(M_n(R)\)

OpenStudy (anonymous):

conversely if \(J\) is an ideal in \(M_n(R)\) then there is some ideal \(I\) in \(R\) with \(J=M_n(I)\)

OpenStudy (anonymous):

that fact that \(\mathbb{R}\)is a field means the only ideal in \(\mathbb{R}\) are \(\mathbb{R}\) and \(0\)

OpenStudy (loser66):

Thank you. Need time to digest it. :)

OpenStudy (anonymous):

yw btw you probably do not get to use the statement i wrote, it is often written as "if \(F\) is simple then \(M_n(F)\) is simple

OpenStudy (loser66):

The terminology "simple" is used for group, right? apply it to ring??

OpenStudy (loser66):

One more question: in book it says Set of unit \(E_{ij}\) is a basis for the linear space of matrices. \(E_{ij}E_{kl}=\delta_{jk}E_{il}\) I don't understand it. Please, explain I know \(E_{ij}\) is a matrix with 1 in the (i,j) position and 0 elsewhere. But don't ge the product above

OpenStudy (loser66):

Let say, in 2x2 matrices, \(E_{1,1}=\left[\begin{matrix}1&0\\0&0\end{matrix}\right]\) is one of the basis. And \(E_{2,1}=\left[\begin{matrix}0&0\\1&0\end{matrix}\right]\) \(E_{1,1}E_{2,1}=0\) , so, what is \(\delta_{jk}\) ? What is \(E_{il}\)?

OpenStudy (anonymous):

i guess this proof is going to go as follows if there is a non zero element in any matrix in the ideal, then the ideal contains all matrices

OpenStudy (anonymous):

so pick some matrix in the ideal that is not identically zero then since that matrix times an other matrix must also be in the ideal, so that you can put an entry anywhere you like in the product of two matrices, where one has one non zero entry

OpenStudy (anonymous):

*show, not "so"

OpenStudy (loser66):

Down the case to 2x2 only, show me,please

OpenStudy (anonymous):

\[\delta_{ij}\] is the dirichlet delta, a fancy way of saying \(1\) if \(i=j, 0\) otherwise

OpenStudy (anonymous):

suppose you are working in 2 by 2 matrices, and suppose you have an ideal that is not identically zero that means there is a matrix in the ideal that contains a non zero element call it what you like

OpenStudy (loser66):

oh, got you, so \(\delta_{jk}= \delta_{1,2}=0\) and \(E_{il}=E_{1,1}\) right?

OpenStudy (anonymous):

say \[\left[\begin{matrix}a&0\\0&0\end{matrix}\right]\] is in the ideal then since the product of any matrix and this matrix is in the ideal, show that the ideal contains all matrices

OpenStudy (anonymous):

i.e you can put any number anywhere you like and don't forget it is closed under addition, so you really just need to show you can put an entry in any position

OpenStudy (loser66):

In matrix, we don't have commutative property, so no matter what happen, matrix can't have any ideal which defined by left ideal = right ideal

OpenStudy (anonymous):

you are supposed to sh ow it can't have any ideal at all!

OpenStudy (loser66):

oh, got it. hehehe...

OpenStudy (loser66):

Thank you so much.

OpenStudy (anonymous):

yw